OpenStudy (anonymous):

N is a set of numbers whose average (arithmetic mean) is 3. M is a set that is generated by doubling each number in N. What is the average of the numbers in set M? 1/3 2 3 6 9

OpenStudy (shane_b):

If you double the individual values the average result will double.

OpenStudy (chihiroasleaf):

Let \[x\] be the sum of numbers in N, then the sum of numbers in M is \[2x\] The number of numbers in N and M are the same, so if \[n\] represents the numbers in N and M, then the average of M is \[\frac{ 2x }{ n }=2\times \frac{ x }{ n }=2\times3=6\]

OpenStudy (shane_b):

If you consider the fact that you're just doubling the numerator...you'll see that you're doubling the average as well.

OpenStudy (anonymous):

so is it 6?

OpenStudy (shane_b):

Yes...